I need help on Expressions involving variables, substitution.
"John paid $34 for two algebra and three geometry books. He paid $36 for three algebra and two geometry books. What is the cost of each book?"
Thank You!

Answer by mananth(16949)

Algebra book cost x
geometry book cost y
2x+3y=34
3y=-2x+34
y=(-2/3)x+34/3
3x+2y=36
substitute y in the equation
3x+2((-2/3)x+34/3)=36
3x-(4/3) x +68/3=36
multiply equation by 3
9x-4x+68=108
5x=40
x=8 , algebra book cost
plug x in equation 2x+3y=34 to get y=6